each side of a square is lengthened by 1 inch. the area of this new, larger square is 100 square inches. find the length of a side of the original square.
Answer by checkley79(3341)

AREA=S^2
AREA=(S+1)^2
100=S^2+2S+1
S^2+2S+1-100=0
S^2+2S-99=0
(S+11)(S-9)=0
S-9=0
S=9 IN. ANS. FOR THE ORIGINAL SIDES.
PROOF:
100=(9+1)^2
100=10^2
100=100